OpenStudy (anonymous):

find the perimeter of a rectangular farm that is 1/2 mile wide and 3/4 mile long

OpenStudy (anonymous):

P = 2(L+B)

OpenStudy (lgbasallote):

\[P = l+ l + w + w\] \[\implies \frac 12 + \frac 12 + \frac 34 + \frac 34\] does that help?

OpenStudy (anonymous):

oh...sorry

OpenStudy (anonymous):

mistake in calculation

OpenStudy (anonymous):

@Igbasallote i came up with 2.5......

OpenStudy (anonymous):

= 2 *(1/2 + 3/4) = 2 * 5/4 = 5/2 = 2.5

OpenStudy (anonymous):

yup u r correct

OpenStudy (lgbasallote):

1/2 + 1/2 + 3/4 + 3/4 = 1 + 6/4 = 1 + 3/2 = 1 + 1/5 = 2.5 right

OpenStudy (anonymous):

cool. thank you

OpenStudy (lgbasallote):

that should be 1 + 1.5 not 1/5